OpenStudy (anonymous):

to determine the order of a reaction do we multiply the intial reaction of first value with the second one in a given data?

OpenStudy (xishem):

How to determine the order of reaction in respect to each reactant species: 1 0.10 M 0.10 M 1.23 x 10-3 M/s 2 0.10 M 0.20 M 2.46 x 10-3 M/s 3 0.20 M 0.10 M 4.92 x 10-3 M/s 4 ? 0.30 M 8.00 x 10-3 M/s

OpenStudy (anonymous):

thanx

OpenStudy (xishem):

Let's first write the general rate equation:\[Rate=k[A]^m[B]^n\]Let's ignore k for now. Since it's constant, it won't affect rates based on concentrations of reactants.\[Rate=[A]^m[B]^n\]In the case of the second row of the table, [A] is constant with respect to the first row, so we can ignore that too. We are doubling [B] and by doing that, the rate is doubled, so:\[2(Rate)=(2[B])^n \rightarrow 2(Rate)=2^n[B]^n\]Rate and [B] are constant for this particular reaction, so we can drop them off:\[2=2^n\]We can see that n=1 here. Now let's look at the 1st and 3rd rows and go through a similar process. k and [B] are constant with respect to the two rows, so:\[Rate=[A]^m\]When the concentration is doubled, the rate is quadrupled:\[\frac{4.92 \times 10^{-3}}{1.23 \times 10^{-3}}=4\]So we can write the rate equation as:\[4(Rate)=(2[A])^m \rightarrow 4(Rate)=2^m[A]\]Since Rate and [A] are constant for this particular case, we can drop them off and we are left with:\[4=2^m\]Which leaves "m" to be what?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so to find the order of a reaction in a given data we have to choose any two experiments?

OpenStudy (xishem):

Yes, and then the final rate reaction equation is:\[Rate=k[A]^2[B]\]

OpenStudy (anonymous):

after taking any two rows we use this formula

OpenStudy (xishem):

Yes, you need to choose two experiments where one of the concentrations is constant.

OpenStudy (anonymous):

or equation rate =k [a][b]

OpenStudy (xishem):

For example, you can't determine the powers if you have, for instance, these two sets of data: 1 0.10 M 0.10 M 1.23 x 10-3 M/s 2 0.20 M 0.20 M 9.84 x 10-3 M/s In this case, it is impossible to determine how the concentrations of each reactant affect the rate because there are no constants to compare to.

OpenStudy (xishem):

You'll always compare two experiments where at least one variable is constant between the two, whether it's [A], [B], or rate.

OpenStudy (anonymous):

oh so in ur given example which is constant?

OpenStudy (xishem):

No variables are constant. The orders of reaction can't be determined by just that data.

OpenStudy (anonymous):

ahan

OpenStudy (anonymous):

my teach used the log method to find the order

OpenStudy (xishem):

If you use the previous method described:\[Rate=k[A]^m[B]^n\]k is constant, so we drop it.\[Rate=[A]^m[B]^n\]If we double both concentrations, we see that the rate is multiplied by 8:\[8(Rate)=(2[A])^m(2[B])^n \rightarrow 8(Rate)=2^m2^n[A]^m[B]^n\]Drop off the variables to get:\[8=2^m2^n\]To which there are an infinite number of solutions.

OpenStudy (anonymous):

thats easier than the log method

OpenStudy (anonymous):

so this is the order of the reaction.

OpenStudy (xishem):

Yeah, it can be easier. The log method is basically a shortcut of doing the calculation once you understand the concept behind it.

OpenStudy (anonymous):

now how can we find the overall order?

OpenStudy (anonymous):

becuase we found the order of the reaction not each reactant

OpenStudy (xishem):

Well, we found this to be true, right?\[Rate=k[A]^2[B]\]

OpenStudy (anonymous):

ohh so we found the order of A

OpenStudy (xishem):

And the order of B. It was just implied as 1:\[Rate=k[A]^2[B]^1\]

OpenStudy (anonymous):

2+1 = 3 overall order is 3

OpenStudy (xishem):

Yes.

OpenStudy (anonymous):

so every time we use log or any method to find the order we are actually finding the order of A only.

OpenStudy (anonymous):

how B is 1? same method?

OpenStudy (xishem):

Yes. Same method. We look at the original table: 1 0.10 M 0.10 M 1.23 x 10-3 M/s 2 0.10 M 0.20 M 2.46 x 10-3 M/s 3 0.20 M 0.10 M 4.92 x 10-3 M/s 4 ? 0.30 M 8.00 x 10-3 M/s And we choose the two rows in which B changes, so we can see how the rate varies with B. These are rows 1 and 2. The general rate equation:\[Rate=k[A]^m[B]^n\]Drop off k and [A]^m because they are constant between rows 1 and 2:\[Rate=[B]^n\]When we double [B], we see that the rate doubles as well, so:\[2(Rate)=(2[B])^n \rightarrow 2(Rate)=2^n[B]^n\]Drop Rate and [B]^n:\[2=2^n \rightarrow n=1\]

OpenStudy (anonymous):

oh so in this reaction we did not find the value of A because it was constant

OpenStudy (anonymous):

we only find the order of B becuase it was not constant?

OpenStudy (anonymous):

I mean the order of A.

OpenStudy (anonymous):

so the order which we find out , is it the order of B ?

OpenStudy (xishem):

Yes. It is impossible to determine the order of A only using rows 1 and 2. We were able to find the order of [B] by knowing how rate changes with [B].

OpenStudy (anonymous):

becuase A was constant then we only check the order of B. But in case if we take the 3 and 1 in which the rate is changing in A then what ?

OpenStudy (xishem):

It's what I outlined before. The third post in the question. My second post.

OpenStudy (anonymous):

so even if I take row 1 or 2 or row 3 and 1. I will come up with the same ordeR?

OpenStudy (xishem):

Yes. As long as you take into account which values are constant when you're working with your equations.

OpenStudy (anonymous):

ok. U mean between A or B atleast one value has to be constant

OpenStudy (xishem):

Yes. Otherwise you end up with an equation like this (which I posted before), which cannot be solved because there are two unknown variables which are not functions of eachother:\[8=2^m2^n\]

OpenStudy (anonymous):

ook . u can see that there is a question mark in the table. which means the last initial concentration of A is missing ? How will you find that? what equation applies here?

OpenStudy (xishem):

Since we've found the orders of A and B, we know that:\[Rate=k[A]^2[B]\]We can choose a specific row on the table where we know rate, [A], and [B] and that leaves only k to be solved for. So go ahead and solve for k by choosing any row -- it doesn't matter which one you choose. They will all give the same answer.

OpenStudy (anonymous):

I choosed 1st row and I came up with 25.92

OpenStudy (anonymous):

into 10^-2

OpenStudy (anonymous):

I applied the equation k = rate/[a][b]

OpenStudy (xishem):

I get 1.23(M^-2)(s^-1)

OpenStudy (anonymous):

ohh yes yes . sorry .

OpenStudy (xishem):

\[Rate=k[A]^2[B] \rightarrow k=\frac{Rate}{[A]^2[B]} \rightarrow k=\frac{1.23 \times 10^{-3}\frac{M}{s}}{(0.10M)^2(0.10M)}=1.23\frac{1}{M^2s}=1.23M^{-2}s^{-1}\]

OpenStudy (anonymous):

yes yes.

OpenStudy (anonymous):

now how will you find the intial rate of A which is missing in the table

OpenStudy (xishem):

Back to the rate equation:\[Rate=k[A]^2[B]\]From the 4th row of the table, we know everything except [A], so we can calculate it, yes?

OpenStudy (anonymous):

yes

OpenStudy (xishem):

And we'll get?

OpenStudy (anonymous):

so we can calculate the 8.00/3.0?

OpenStudy (xishem):

\[Rate=k[A]^2[B] \rightarrow [A]=\sqrt{\frac{Rate}{k[B]}}\]We know rate and [B] for row 4 and we just calculated k. Plug the numbers in an go.

OpenStudy (anonymous):

so where the sq root came from?

OpenStudy (xishem):

Algebraic manipulation of the rate equation:\[Rate=k[A]^2[B] \rightarrow \frac{Rate}{k[B]}=[A]^2 \rightarrow \sqrt{\frac{Rate}{k[B]}}=[A]\]Make sense?

OpenStudy (anonymous):

ohh . so I just add the values and that will be my missing rate

OpenStudy (anonymous):

I didnot get it ?

OpenStudy (anonymous):

In the above equation do I have to add any values?

OpenStudy (xishem):

You aren't missing a rate, you are missing a concentration. You don't add any values, you just substitute the values that you know from row 4 of the table, which are rate, [B], and we've calculated k above. Plug those numbers in and solve for [A].

OpenStudy (anonymous):

rate of the same row in which the A is missing?

OpenStudy (xishem):

Yes.

OpenStudy (anonymous):

yes I found it .it is 0.215

OpenStudy (anonymous):

0.021

OpenStudy (anonymous):

am I right?

OpenStudy (xishem):

I'm getting 0.147M.

OpenStudy (anonymous):

do I ALSO have to take under root of 0.021

OpenStudy (xishem):

Yes. 0.021 is [A]^2

OpenStudy (anonymous):

ok .. Thanx alot u taught me a great concept.

OpenStudy (anonymous):

how will you find the rate if its missing?

OpenStudy (xishem):

You would need to be given the concentrations of both reactants, and you could just use the rate reaction equation in its standard form. Plug in k and both concentrations to get rate:\[Rate=k[A]^2[B]\]

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

when we calculate the log why it came up with 0?

OpenStudy (xishem):

The log of what?

OpenStudy (anonymous):